OpenStudy (anonymous):

when you multiply it all out, how many digits are in the answer to 2^350 x 3^3 x 5^351

OpenStudy (anonymous):

2^350 * 5^351 * 3^3 = 2^350 * 5^350 * 5 * 27 = (2*5)^350 * 135 = 135 * 10^350 = 1.35 * 10^352 so, the answer has 352 digits.

OpenStudy (anonymous):

@fiddlearound you are off by a digit

OpenStudy (anonymous):

yesss :) lucky the question was posted twice !!

OpenStudy (anonymous):

\[3.14\times 10^3=3140\] has 4 digits